# TimeLDM: Latent Diffusion Model for Unconditional Time Series Generation

### v1

@@ 5 Conclusion
In this paper, we propose TimeLDM, a novel latent diffusion model for unconditional time series generation. Particularly, we explore diffusion on the informative and smoother latent space where the original time series is encoded by a variational autoencoder. We evaluate our method on the simulated and realistic datasets, benchmark the performance against existing state-of-the-art methods. Experimental results demonstrate that TimeLDM persistently delivers high-quality generated data both qualitatively and quantitatively. Scores from Context-FID and Discriminative indicate that TimeLDM surpasses the current state-of-the-art in time-series data generation with an average improvement of 3.4$\times$ and 3.8$\times$, respectively. Further studies demonstrate that our method yields better performance on different lengths of time series data generation. To the best of our knowledge, this is the first work to explore the potential of the latent diffusion model for unconditional time series generation.
{ack}
We would like to thank all the anonymous reviewers for their constructive feedback.

### v2

@@ V CONCLUSIONS
In this paper, we propose TimeLDM, a novel latent diffusion model for unconditional time series generation. Particularly, we explore diffusion on the latent space where the original time series is encoded by a variational autoencoder. We evaluate our method on the simulated and real-world datasets and benchmark the performance against existing state-of-the-art methods. Experimental results demonstrate that TimeLDM persistently delivers high-quality generated data both qualitatively and quantitatively. Remarkably, TimeLDM achieves new state-of-the-art results on the simulated benchmarks and an average improvement of 55% in Discriminative score with all benchmarks. Further studies demonstrate that our method yields better performance on different lengths of time series data generation. To the best of our knowledge, this is the first work to explore the potential of the latent diffusion model for unconditional time series generation. We hope that TimeLDM can serve as a robust baseline for generating informative time series tokens for agents learning in the field of physical AI.